Question 1 18 pts 1) A charge, q1, of -9 micro-coulombs is located at the origin, a charge, q2, of -10 micro-coulombs is located at x = 0 cm, y = +9.9 cm, a charge, q3, of +8 micro- coulombs is located at x 0 cm, y #-3.2 cm, and a charge, q4, of +4 micro- coulombs is located at x = 9.7 cm, y = 0 cm, what is the magnitude of the total electric force on q2 (charge on the ty axis) in Newtons?
Answer #1

Force diagram for the given problem is shown below ;

(o,9.9) A4 、 (97 , 0) (0 ,-3.2) C

Where, F_{4} is force on the charge q_{2} due to the charge q_{4}

  F_{3}  is force on the charge q_{2} due to the charge q_{3}

F_{1} is force on the charge q_{2} due to the charge q_{1}

Given, q_{1} = - 9 x 10-6 C located at origin

q_{2} = -10 x 10-6 C located at A (0 , 9.9 cm) or A (0 , 9.9 x 10-2 m)

q_{3} = 8 x 10-6 C located at C (0 , -3.2 cm) or C (0 , -3.2 x 10-2 m)

  q_{4} = 4 x 10-6 C located at B (9.7 cm , 0) or B (9.7 x 10-2 m , 0)

The electrostatic force between two charge particles q and Q and separated by a distance r is given by

1 qe

where,   9 × 109 N7722/C2 4TTEQ

Using this formula we can calculate the magnitude of force acting on q_{2} due to all other particles. (here i am only calculating the magnitude of forces. So, i will not consider the polarity of charges. The direction of force is already shown in the diagram).

(9 × 10-o) × (10 × 10-0) OA2 q192 1

where, OA = 9.9 x 10-2 m

So,

In vector notation :

F1 = 82.64 j (in N)

and

9342 1 (8 × 10-0) (10 x 10-0) F3 (OA +OC)2

Now, OA + OC = (9.9 + 3.2) x 10-2 m = 13.1 x 10-2 m

So, (8 × 10-6) × (10 × 10-6) (13.1 x 10-2)2 F3-9×109 × 41.95 V   

In vector notation

  vec{F}_{3}= 41.95 , , (-hat{j}) , , (in , N)

And finally

(4 × 10-6) × (10 × 10-6) (AB)2 9492 1

In the right angle triangle AOB , AB is the hypotenuse of the triangle.

So,,   AB = sqrt{AO^{2} + OB^{2}} = sqrt{(9.9 imes 10^{-2})^{2} + (9.7 imes 10^{-2})^{2}}

V 192.1 × 10-4 13.86× 10-2 m 7n

So,   9 .. (4 × 10-6) × (10 × 10-6) F4-9×109 × = 18.74 N (13.86 x 10-2)2

And, from the figure

  OB 9.7 tare = ー 0.98 _ 0.98 _ Q4 9.9

or,  θ tan-10.98 44.415°

So, in vector notation F_{4} can be written as

vec{F}_{4} = F_{4} , cos heta , , (- hat{j}) + F_{4} , sin heta , , hat{i}

or,   F,--18.74 × cos 44.415°-j + 18.74 × sin 44.415%

or, vec{F}_{4} =-13.386 , , hat{j} + 13.12 , , hat{i} (in N)

So, total electric force on q_{2} can be given as

vec{F}= vec{F}_{1} + vec{F}_{3} + vec{F}_{4}

or, vec{F}= left ( 82.64 , , hat{j} ight ) + left ( -41.95 , , hat{j} ight ) + left ( -13.386 , , hat{j} + 13.12 , , hat{i} ight )

or, vec{F}= 27.304 , , hat{j} + 13.12 , , hat{i}

So, magnitude of total electric force on q_{2} is

F = | F| = ν/(27.304 )2 + (13.12)2 = V9 17.643 = 30.3 V

For any doubt please comment and please give an up vote. Thank you.